OpenStudy (chycora):

Molly's alarm clock fell to the floor and the glass face was broken. Molly measured the glass face so that she could order a new one. The radius of the circle-shaped glass face is 6 centimeters. What is the approximate area of the glass face? (Use 3.14 for .) 113.04 cm2 37.68 cm2 452.16 cm2 18.84 cm2

Mehek (mehek14):

do you know the formula

OpenStudy (anonymous):

As i said earlier \[A = 2 \times \pi \times r^2\]

OpenStudy (anonymous):

where pi = 3.14 r = 6

OpenStudy (anonymous):

whats is 6^2?@chycora ?

OpenStudy (anonymous):

in other words what is 6 times 6

OpenStudy (chycora):

it's 36

OpenStudy (anonymous):

Good, so A = 2 times 3.14 times 36

OpenStudy (chycora):

soA = 2 times 3.14 times 36=226.08

OpenStudy (anonymous):

Yes!

OpenStudy (chycora):

ok wat next cuz that's nt the answer

OpenStudy (anonymous):

A

OpenStudy (chycora):

ok thanx